OpenStudy (anonymous):

9. What is the measure of one of the interior angles in a regular convex hexagon?

OpenStudy (anonymous):

120

OpenStudy (anonymous):

chris - mathlearner is wrong on 15 polygon

OpenStudy (anonymous):

r u shure

OpenStudy (anonymous):

absolutely certain the total is 15 * 156 = 2340 his formula is wrong check it out on google

OpenStudy (anonymous):

using the formula for calculating total angles in 15-gon = (n-2) * 180 = 13 * 180 = 2340 same as i got by the other method